calculate the energy stored in a inductor of inductance 100mH when a current of 4.0 A is passed through it?

Grade:

1 Answers

Nirmal Singh.
9 years ago
Energy stored in a inductor

E = 1/2 L*i^2

E= 0.5*100*10^-3*(4)^2

E = 0.5*0.1*16
E = 0.80 J

Regards,
Nirmal SIngh
Askiitians Faculty